OpenStudy (anonymous):

find the exact solution, using common logarithms, and a two-decimal-place approximation of the solution 4^x-3(4^-x)=8

OpenStudy (agent0smith):

\[\Large 4^x-3(4^{-x})=8\]first subtract 8 on both sides \[\Large 4^x-3(4^{-x})-8=0\]I would multiply both sides by 4^x \[\Large 4^{2x}-3-8(4^x) = 0\](think about exponent laws and you'll see how to get it. Now it's like a quadratic.

OpenStudy (anonymous):

im terribly confused as to how i would apply common logarithms to this problem. :(

OpenStudy (mertsj):

You could say: Let 4^x=y then you would have : y^2-8y-3=0 Solve for y using the quadratic formula. Set each answer equal to 4^x and take the log of both sides.

OpenStudy (anonymous):

 i ended up with this